OpenStudy (anonymous):

If f(x)=e^sinx, how many zeros does f'(x) have on the closed interval [0,2╥]?

OpenStudy (amistre64):

cos(x) * e^(sin(x)) = 0 when x = pi/2 and 3pi/2

OpenStudy (anonymous):

e^(sin[x])*cos[x]=0 and solve for x it is zero when cos[x] is zero namely; at pi/2, 3pi/2 in [0,2pi]

OpenStudy (anonymous):

How'd you guys get cos(x)??

OpenStudy (amistre64):

its the chain rule of derivatives

OpenStudy (amistre64):

Du e^u; u = sin(x) cos(x) e^(sin(x))

OpenStudy (anonymous):

got it! Thanks :)